Two equal amounts of money were invested in two different stocks. The value of the first stock increased by 15% the first year and decreased by 15% the second year. The second stock decreased by 15% the first year and increased by 15% the second year. What investment was more profitable?
Answer by ikleyn(52781)

.
Let P be the equal amount invested initially in each of the two stocks.


Then the first amount became  P*1.15*0.85 = 0.9775*P at the end of the second year.


     The second amount became  P*0.85*1.15 = 0.9775*P at the end of the second year.


As you see, both investments were equally profitable with the negative profit (= loss) of  (100% - 97.75%) = 2.25%  at the end of the second year.